# Stationary densities and delocalized domain walls in asymmetric exclusion processes competing for finite pools of resources

### v1

@@ V.1 LD-LD phase
Inequalities (20), (23), (24) together define the LD-LD region in the $\alpha$-$\beta$ plane for a fixed $\mu$, see the LD phase regions in the phase diagrams of Figs. 2 and 8.
@@
At this point, we find the range of $\mu$ that permits LD-LD phase in the phase diagram. Let $\mu=3/2-\epsilon$, with $0<\epsilon<3/2$. Substituting this into inequality (20) gives
$\alpha>\frac{1}{2}-\epsilon,$ (26)
and inequality (24) becomes
$\frac{1}{\alpha}>2(1-\epsilon).$ (27)
For $0<\epsilon<1$, inequality (27) implies
$\alpha<\frac{1}{2(1-\epsilon)},$ (28)
which, when combined with (26), yields a consistent range for $\alpha$. For $1<\epsilon<3/2$, inequality (27) instead gives
$\alpha>2(1-\epsilon),$ (29)
which is again compatible with (26). Thus, for all $0<\epsilon<3/2$, i.e., for $0<\mu<3/2$, the inequalities (20), (23), and (24) defining the LD-LD phase region are simultaneously satisfied. Now consider $\mu=3/2+\epsilon$, with $0<\epsilon<1/2$. Substituting this into (20) gives
$\alpha>\frac{1}{2}+\epsilon,$ (30)
while (24) translates
$\alpha<\frac{1}{2(1+\epsilon)}.$ (31)
However, since
$\frac{1}{2(1+\epsilon)}<\frac{1}{2}<\frac{1}{2}+\epsilon$
for $0<\epsilon<1/2$, the two inequalities (30) and (31) are incompatible. This contradiction shows that the LD-LD phase cannot exist for $3/2<\mu<2$.
This is consistent with the Monte Carlo simulation (MCS) results in Figs. 2 and 8 where the LD-LD phase appears for
$0<\mu<\frac{3}{2},$ (32)
and is absent for $3/2<\mu<2$.
@@ V.2 HD-HD phase
This phase also corresponds to the solution $\rho^{(1)}=\rho^{(2)}\equiv\rho_{\text{HD-HD}}>1/2$. In the HD-HD phase, characterized by high densities in both TASEP lanes, the bulk densities are given by $\rho^{(1)}=1-\beta_{\text{eff}}^{(1)}>1/2$ and $\rho^{(2)}=1-\beta_{\text{eff}}^{(2)}>1/2$. Imposing $\rho^{(1)}=\rho^{(2)}$ leads to
$1-\beta_{\text{eff}}^{(1)}=1-\beta_{\text{eff}}^{(2)}$
$\implies\quad$ $N_{1}=N_{2}.$ (33)
@@
Again, $N_{1}$ and $N_{2}$ scale linearly with the system size $L$, and depend linearly on the filling factor $\mu$. Since $0<N_{1},N_{2}\leq L$, Eq. (34) yields
$\frac{2-\mu}{1+\beta}<1\quad\text{and}\quad\mu\leq 2.$ (35)
@@
Inequalities (35), (38), (39) together define the HD-HD region in the $\alpha$-$\beta$ plane for a fixed $\mu$, see the HD region displayed in the phase diagrams of Figs. 2 and 8.
@@
The range of $\mu$ for HD-HD phase to exist is obtained from MFT as following. Assume $\mu=1/2+\epsilon$, where $0<\epsilon<3/2$. Substituting into inequality (35) yields:
$\beta>\frac{1}{2}-\epsilon,$ (41)
while inequality (39) becomes:
$\frac{1}{\beta}>2(1-\epsilon).$ (42)
For $0<\epsilon<1$, inequality (42) implies:
$\beta<\frac{1}{2(1-\epsilon)},$ (43)
which, when combined with (41), provides a consistent range for $\beta$, confirming the feasibility of the HD-HD phase in this regime. For $1<\epsilon<3/2$, inequality (42) instead implies:
$\beta>2(1-\epsilon),$ (44)
which also remains compatible with (41). Therefore, the inequalities (35), (38), and (39) are simultaneously satisfied for the full range $0<\epsilon<3/2$, i.e., for $1/2<\mu<2$, justifying the existence of the HD-HD phase in this regime. We now show that the HD-HD phase does not occur for $0<\mu<1/2$. Let $\mu=1/2-\epsilon$, with $0<\epsilon<1/2$. Substituting this into inequality (35) gives:
$\beta>\frac{1}{2}+\epsilon,$ (45)
while (39) becomes
$\beta<\frac{1}{2(1+\epsilon)}.$ (46)
However, for $0<\epsilon<1/2$, we have:
$\frac{1}{2(1+\epsilon)}<\frac{1}{2}<\frac{1}{2}+\epsilon,$
so the inequalities (45) and (46) are mutually incompatible. This contradiction excludes the possibility of the HD-HD phase when $0<\mu<1/2$.
These analytical results are consistent with the MCS results presented in the phase diagrams of Fig. 2 and Fig. 8 where the HD-HD phase is found for
$\frac{1}{2}<\mu<2,$ (47)
and is absent for $0<\mu<\frac{1}{2}$.
Figure 3: (a) Plot of $\rho_{\text{LD}}$ vs $\alpha$ for a fixed $\mu$, i.e., $\mu=0.3$, see Eq. (25). For large values of $\alpha$, $\rho_{\text{LD}}\rightarrow\mu$. (b) Plot of $\rho_{\text{HD-HD}}$ vs $\beta$ for a fixed $\mu$, i.e., $\mu=1.7$, see Eq. (40). For large values of $\beta$, $\rho_{\text{HD-HD}}\rightarrow\mu-1$. MFT and MCS results are in great agreement.
@@ V.3 MC-MC phase
in the MC-MC phase. These imply
$\frac{1}{2\alpha}<\frac{N_{i}}{L}<1-\frac{1}{2\beta},\quad i=1,2,$ (52)
@@
which again cannot be satisfied for $\beta>0$. This rules out the MC-MC phase for $\mu>3/2$. Therefore, the MC-MC phase is restricted to the interval $1/2<\mu<3/2$, in full agreement with the MCS observations in Figs. 2 and 8.
This result can also be understood from particle number conservation. In the MC-MC phase, each lane contains $L/2$ particles, so the total number of particles in the lanes is $L$. Using the relation $N_{0}=2\mu L$, particle number conservation implies:
@@ V.4 DW-DW phase
where $\Theta(x)$ is the Heaviside step function defined as $\Theta(x)=1(0)$ for $x>(<)0$. Similar to an open TASEP, one has $\rho_{\text{LD}}^{(i)}=\alpha_{\text{eff}}^{(i)}$ and $\rho_{\text{HD}}^{(i)}=1-\beta_{\text{eff}}^{(i)}$ in the DW-DW phase of the present model. Now, current conservation in the DW-DW phase implies $\rho_{\text{LD}}^{(i)}+\rho_{\text{HD}}^{(i)}=1$ which, along with the solution $\rho^{(1)}=\rho^{(2)}$, infers $\alpha_{\text{eff}}^{(1)}=\alpha_{\text{eff}}^{(2)}=\beta_{\text{eff}}^{(1)}=\beta_{\text{eff}}^{(2)}$. This in turn yields:
$\frac{N_{1}}{L}=\frac{\beta}{\alpha+\beta}=\frac{N_{2}}{L}$ (62)
in the DW-DW phase. Thus, effective entry and exit rates [see Eqs. (2)-(4)] in both lanes are obtained as
@@
This equation cannot be solved to determine both $x_{w}^{(1)}$ and $x_{w}^{(2)}$ uniquely keeping the parameters $\alpha,\,\beta$ and $\mu$ fixed, thereupon, indicating the domain walls to be delocalized in space. Physically, shifting a DW in any direction in one lane is compensated by an opposite shift of the DW in another lane, keeping PNC unaffected. Thus all such pairs of $(x_{w}^{(1)},x_{w}^{(2)})$ give valid DW positions. The inhererent stochasticity of the dynamics ensures that eventually the system visits all such solutions, displaying only the envelop of all the DWs under long time averaging. Since each channel is uniform having unit hopping rate in the bulk, the probability of finding a DW in a given channel is a constant. This in turn gives that the envelop of such a DW as an inclined lines, i.e., a DDW profile, spanning either the full length of the each lane, or only partially. In case of a partially spanning DDW, the fraction spanned by a DDW in each of the lanes is equal. This is a consequence of PNC. Thus in the DW-DW phase, the domain walls in both lanes become delocalized due to the freedom in their relative positions, constrained only by the total particle number conservation, see Eq. (67), see also Ref. sm-inhomogeneous-tasep .
@@
Inequalities (65) and (69) together define the DW-DW region in the $\alpha$-$\beta$ phase diagram for any $0\leq\mu\leq 2$. This concludes the DW-DW phase is always present in our model, see Figs. 2(a)-2(d) and Figs. 8(a), 8(b).
We now examine whether the DDWs are completely or partially delocalized. The condition $\rho^{(1)}=\rho^{(2)}$ along with PNC, implies that the time-averaged positions of both domain walls satisfy $\langle x_{w}^{(1)}\rangle=\langle x_{w}^{(2)}\rangle=x_{0}$ in the steady state, where $\langle\cdot\rangle$ denotes an average over steady-state realizations. In a closed system with biased hopping, particles tend to accumulate behind local inhomogeneities (in this case, the reservoirs). Assuming linear profiles for the DDWs, the average domain wall position $x_{0}$ can be determined from PNC as:
@@
Equation (71) shows, for fixed values of $\alpha$ and $\beta$, $x_{0}$ decreases with increasing $\mu$, indicating that the domain walls shift toward the entry-end of the respective lane.
Table 1 analyzes the DDWs using a geometric approach, summarizing the span $\Delta$ corresponding to different values of $x_{0}$ and the associated conditions on $\mu$. The geometric constructions of DDWs, illustrated in Fig. 6, show excellent agreement with MCS results.
@@
Figure 4: Plots showing the variation of system size-averaged density with Monte Carlo sweeps for (a) fully delocalized domain wall with parameters $\alpha=\beta=0.5,\mu=1$ and (b) partially delocalized DW with parameters $\alpha=0.2,\beta=0.5,\mu=1$. System size is $L=1000$ and $3\times 10^{7}$ MC steps are taken. This shows that fluctuations in stationary densities survive even at thermodynamic limit in the DW-DW phase of our model.
Figure 5: (a) Variance of the system-averaged density, $\rho$, versus system size $L$ for a fully delocalized domain wall with $\alpha=\beta=0.5$ and $\mu=1$. Clearly, $\mathrm{Var}(\rho)$ is of order 1. (b) Log–log plot of $\mathrm{Var}(\rho)$ versus $L$ for the LD-LD phase with $\alpha=0.5$, $\beta=1.5$, and $\mu=1$, showing that $\mathrm{Var}(\rho)\sim O(1/L)$.
@@ VI Phase boundaries and phase transitions
The phase diagrams of our model (Figs. 2 and 8) differ substantially from that of the open TASEP, or from other models of TASEPs with finite resources. We identify the boundaries delineating different phases according to MFT. This can be achieved by equating the currents in corresponding phases.
@@
Boundaries (83) and (84) appear as straight lines parallel to the $\beta$ and $\alpha$ axes, respectively, see the phase diagrams in Figs. 2(b)-2(d) and Figs. 8(a)-8(b). The requirement $\alpha,\beta>0$ in Eqs. (83) and (84) imposes the following range on $\mu$ for which these boundaries can exist:
@@
Depending on the value of $\mu$, these boundaries can be either straight lines or curved ones, as shown in Figs. 2 and 8.
We now discuss the nature of the phase transitions across these phase boundaries. The $\alpha$$\beta$ phase space of our model can exhibit simultaneous coexistence of either two or four phases, depending sensitively on the value of $0\leq\mu\leq 2$. For $0<\mu<1/2$, only the LD-LD and DW-DW phases are present; while for $3/2<\mu<2$, only the HD-HD and DW-DW phases are observed. In contrast, for intermediate values ($1/2<\mu<3/2$), all four phases – LD-LD, HD-HD, MC-MC, and DW-DW – coexist and meet at a multicritical point (see Eq. (88)). Due to particle-hole symmetry, the phase diagram for $\mu>1$ can be obtained from that for $\mu<1$. Transitions between different phases – for example, from LD-LD or HD-HD to MC-MC or DW-DW – are reflected as continuous transition in the bulk density, indicating that these are continuous transitions.
Interestingly, Figs. 2 and 8 reveal that, for $1/2<\mu<3/2$, all four phase boundaries intersect at a common point. The transitions across this common meeting point between any two phases are also continuous in nature, since the phase boundaries that meet at this common meeting point are all continuous transition boundaries. Thus this is a multicritical point. The coordinates of this multicritical point in the $\alpha$-$\beta$ phase diagram can be obtained by simultaneously solving the boundary equations (83), (84), (86), and (87):
@@
As $\mu$ varies, the multicritical point ($8$) traces a hyperbolic trajectory (cf. Fig. 7(a)):
@@
Notably, $d$ diverges as $\mu\rightarrow(1/2)^{+}$ and $\mu\rightarrow(3/2)^{-}$, indicating that the multicritical point exists only within the range $1/2<\mu<3/2$ (cf. Fig. 7(b)). The symmetric nature of the plot of $d$ versus $\mu$ about $\mu=1$ reflects the particle-hole symmetry of the model.

### v2

@@ V.1 LD-LD phase
Inequalities (20), (23), (24) together define the LD-LD region in the $\alpha$-$\beta$ plane for a fixed $\mu$, see the LD phase regions in the phase diagrams of Figs. 2 and 10.
@@
At this point, we find the range of $\mu$ that permits LD-LD phase in the phase diagram. This follows from the inequalities (20) and (24), which respectively provide the lower and upper bounds on $\alpha$. These bounds can be satisfied simultaneously only when the parameter $\mu$ lies in the range
$0<\mu<\frac{3}{2}.$ (26)
This is consistent with the Monte Carlo simulation (MCS) results in Figs. 2 and 10 where the LD-LD phase appears for $0<\mu<3/2$ and is absent for $3/2<\mu<2$.
@@ V.2 HD-HD phase
The HD-HD phase is particle-hole symmetric counterpart to the LD-LD phase. Consequently, by applying the transformations (9)–(11), the effective entry and exit rates, phase region and steady state density of the HD-HD phase can be directly obtained from the results of the LD-LD phase discussed in the previous section.
Similar to the LD-LD phase, both reservoirs are equally populated in the HD-HD phase (see Appendix C):
@@
The effective entry and exit rates in the HD-HD phase follow directly from the corresponding effective rates of the LD-LD phase, given in Eqs. (21) and (22), upon applying the particle-hole transformations (9) and (11):
@@
The HD-HD region in the parameter space can similarly be obtained from the LD-LD region, defined by inequalities (20), (23) and (24), using the same set of transformations (9) and (11). This yields the following conditions:
$\mu\leq 2\quad\text{and}\quad\frac{2-\mu}{1+\beta}<1,$ (30)
@@
Inequalities (30), (31), (32) together define the HD-HD region in the $\alpha$-$\beta$ plane for a fixed $\mu$, see the HD region displayed in the phase diagrams of Figs. 2 and 10.
The steady-state densities in both lanes in the HD-HD phase can be obtained from Eq. (25) using the trasformations (10) and (11):
@@
The range of $\mu$ over which the HD-HD phase exists can be directly obtained from the corresponding range for the LD-LD phase by applying the transformation $\mu\to 2-\mu$; see the transformation (11). This yields the following condition for the HD-HD phase:
$\frac{1}{2}<\mu<2.$ (34)
These analytical results are consistent with the MCS results presented in the phase diagrams of Fig. 2 and Fig. 10 where the HD-HD phase is found for $1/2<\mu<2$ and is absent for $0<\mu<\frac{1}{2}$.
Figure 3: (a) Plot of $\rho_{\text{LD-LD}}$ as a function of $\alpha$ for fixed $\mu=0.3$ [see Eq. (25)]. For large $\alpha$, $\rho_{\text{LD-LD}}$ approaches $\mu$. (b) Plot of $\rho_{\text{HD-HD}}$ as a function of $\beta$ for fixed $\mu=1.7$ [see Eq. (33)]. For large $\beta$, $\rho_{\text{HD-HD}}$ approaches $\mu-1$. MFT and MCS results show excellent agreement.
@@ V.3 MC-MC phase
in the MC-MC phase. These give
$\frac{1}{2\alpha}<\frac{N_{a}}{L}<1-\frac{1}{2\beta},\quad a=1,2,$ (39)
@@
which again cannot be satisfied for $\beta>0$. This rules out the MC-MC phase for $\mu>3/2$. Therefore, the MC-MC phase is restricted to the interval $1/2<\mu<3/2$, in full agreement with the MCS observations in Figs. 2 and 10.
This result can also be understood from the particle number conservation. In the MC-MC phase, each lane contains $L/2$ particles, so the total number of particles in the lanes is $L$. Using the relation $N_{0}=2\mu L$, particle number conservation implies:
@@
Notice that in the MC-MC phase, the MFT gives only ranges of $N_{1}$ and $N_{2}$; see Eq. (39) above. We examine this further numerically and solve for $N_{1},\,N_{2}$ numerically in the steady state, averaged over large time. This gives us $\langle N_{1}\rangle$ and $\langle N_{2}\rangle$, where $\langle.\rangle$ implies temporal averages in the steady states. Our numerical results are shown in Fig. 4, which clearly establishes $N_{1}=N_{2}$ within error bars.
Figure 4: Steady state time-averaged reservoir populations $\langle N_{1}\rangle,\langle N_{2}\rangle$ when the TASEP lanes are in their MC-MC phase. Within error bars, $\langle N_{1}\rangle=\langle N_{2}\rangle$ is found.
@@ V.4 DW-DW phase
where $\Theta(x)$ is the Heaviside step function defined as $\Theta(x)=1(0)$ for $x>(<)0$. Similar to an open TASEP, one has $\rho_{\text{LD}}^{(i)}=\alpha_{\text{eff}}^{(i)}$ and $\rho_{\text{HD}}^{(i)}=1-\beta_{\text{eff}}^{(i)}$ in the DW-DW phase of the present model. Now, current conservation in the DW-DW phase implies $\rho_{\text{LD}}^{(i)}+\rho_{\text{HD}}^{(i)}=1$ which, along with the solution $\rho^{(1)}=\rho^{(2)}$, infers $\alpha_{\text{eff}}^{(1)}=\alpha_{\text{eff}}^{(2)}=\beta_{\text{eff}}^{(1)}=\beta_{\text{eff}}^{(2)}$. This in turn gives
$\frac{N_{1}}{L}=\frac{\beta}{\alpha+\beta}=\frac{N_{2}}{L},$ (49)
i.e., $N_{1}=N_{2}$ in the DW-DW phase. Thus, effective entry and exit rates [see Eqs. (2)-(4)] in both lanes are obtained as
@@
This equation cannot be solved to determine both $x_{w}^{(1)}$ and $x_{w}^{(2)}$ uniquely keeping the parameters $\alpha,\,\beta$ and $\mu$ fixed, thereupon, indicating the domain walls to be delocalized in space. Physically, shifting a DW in any direction in one lane is compensated by an opposite shift of the DW in another lane, keeping PNC unaffected. Thus all such pairs of $(x_{w}^{(1)},x_{w}^{(2)})$ give valid DW positions. The inhererent stochasticity of the dynamics ensures that eventually the system visits all such solutions, displaying only the envelop of all the DWs under long time averaging. Since each channel is uniform having unit hopping rate in the bulk, the probability of finding a DW in a given channel is a constant. This in turn gives that the envelop of such a DW as an inclined lines, i.e., a DDW profile, spanning either the full length of the each lane, or only partially. In case of a partially spanning DDW, the fraction spanned by a DDW in each of the lanes is equal. This is a consequence of PNC. Thus in the DW-DW phase, the domain walls in both lanes become delocalized due to the freedom in their relative positions, constrained only by the total particle number conservation, see Eq. (54), see also Ref. [63].
@@
Inequalities (52) and (56) together define the DW-DW region in the $\alpha$-$\beta$ phase diagram for any $0\leq\mu\leq 2$. This concludes the DW-DW phase is always present in our model, see Figs. 2(a)-2(d) and Figs. 10(a), 10(b).
We now examine whether the DDWs are completely or partially delocalized. The condition $\rho^{(1)}=\rho^{(2)}$ along with PNC, implies that the time-averaged positions of both domain walls satisfy $\langle x_{w}^{(1)}\rangle=\langle x_{w}^{(2)}\rangle=x_{0}$ in the steady state, where $\langle\cdot\rangle$ denotes time-average over steady-state realizations. In a closed system with biased hopping, particles tend to accumulate behind local inhomogeneities (in this case, the reservoirs). Assuming linear profiles for the DDWs, the average domain wall position $x_{0}$ can be determined from PNC as:
@@
Equation (58) shows, for fixed values of $\alpha$ and $\beta$, $x_{0}$ decreases with increasing $\mu$, indicating that the domain walls shift toward the entry-end of the respective lane. Next, we determine the span ($\Delta$) of the DDW, defined as the spatial extent over which the time-averaged DDW density profile (appearing as an inclined line) persists. Table 1 analyzes the DDWs using a geometric approach, summarizing the span $\Delta$ corresponding to different values of $x_{0}$ and the associated conditions on $\mu$. The geometric constructions of DDWs, illustrated in Fig. 5, show excellent agreement with MCS results.
@@
To explore the extent of number fluctuations in the DW-DW phase, going beyond the MFT, we study the model numerically by MCS. We have noted that the DDWs fluctuate over a distance that scales with $L$. Therefore, the fluctuations in the total particle numbers in the TASEP lanes also scale with $L$ in the DW-DW phase. See Fig. 6 for plots of the densities in the TASEP lanes, averaged over entire TASEP length, i.e., $L$ as functions of time, measured by the number of Monte-Carlo steps. The plots reveal complementary nature of the fluctuating densities in the two TASEP lanes.
To study these density fluctuations quantitatively, we compute the standard deviations $\sigma_{i}^{\rho}$ of the system size-averaged densities $\bar{\rho}_{i}$ ($i=1,2$ for the two TASEP lanes), obtained from different Monte Carlo steps. They are defined as
$\sigma_{i}^{\rho}=\sqrt{\langle\bar{\rho}_{i}^{2}\rangle-\langle\bar{\rho}_{i}\rangle^{2}},\quad i=1,2,$ (59)
where
$\bar{\rho}_{i}=(\rho_{\text{HD}}^{(i)}-\rho_{\text{LD}}^{(i)})\,\Delta+\rho_{\text{LD}}^{(i)}$ (60)
in the DW-DW phase, $\rho_{\text{LD}}^{(i)}$ and $\rho_{\text{HD}}^{(i)}$ are the stationary densities in the low- and high-density domains of a DW obtained in Eq. (51). The DDW span $\Delta$ is same (as found in MFT above, see Table 1) in both the TASEP lanes. We have calculated $\sigma^{\rho}_{i}$ for various $L$ from our MCS studies, and found them to be independent of $L$ in the DW-DW phase for a range of $L$ used in our simulations; see Fig. 7(a). Since both $\bar{\rho}_{1},\bar{\rho}_{2}$ are $\mathcal{O}(1)$} quantities independent of $L$, this strongly suggests that the relative fluctuations in $\rho_{1},\,\rho_{2}$ about their mean values $\bar{\rho}_{1},\,\bar{\rho}_{2}$ do not vanish in the DW-DW phase in the thermodynamic limit. This is in contrast to the behavior of $\sigma^{\rho}_{i}$ in the LD-LD phase; see Fig. 7(b) with similar behavior expected in the HD-HD and MC-MC phases (not shown). Thus the relative fluctuations in the TASEP densities vanish in the thermodynamic limit in the LD-LD, HD-HD, and MC-MC phases. A pertinent question is whether the reservoir populations $N_{1},\,N_{2}$ also display large fluctuations in the thermodynamic limit. We study this question numerically and the results are shown in Fig. 8(a) for DW-DW phase and in Fig. 8(b) for LD-LD phase, which shows a monotonically decaying trend with a rising $L$, strongly suggesting the relative fluctuations in the reservoir populations vanish in the thermodynamic limit. This together with the overall particle number conservation explains the observed complementary nature of the fluctuations in the two TASEP densities found in Fig. 6. We thus conclude that in the DW-DW phase while the TASEP densities exhibit large fluctuations that survive in the thermodynamic limit, the relative fluctuations in the reservoir populations vanish in the thermodynamic limit.
Figure 6: Plots showing the variation of system size-averaged density $\bar{\rho}_{i}$ ($i=1,2$ for the two TASEP lanes) with Monte Carlo steps for (a) fully delocalized domain walls with parameters $\alpha=\beta=0.5,\mu=1$ and (b) partially delocalized domain walls with parameters $\alpha=0.2,\beta=0.5,\mu=1$. System size is $L=1000$ and $3\times 10^{7}$ MC steps are taken. This shows that fluctuations in stationary densities survive even at thermodynamic limit in the DW-DW phase of our model.
Figure 7: Standard deviations $\sigma^{\rho}_{i}$ [see Eq. (59)] of the system size-averaged densities $\bar{\rho}_{i}$ ($i=1,2$ for TASEP lanes $T_{1}$ and $T_{2}$) as functions of system size $L$: (a) for a fully delocalized domain wall with $\alpha=\beta=0.5$ and $\mu=1$, where $\sigma^{\rho}_{i}\sim\mathcal{O}(1)$, and (b) for the LD–LD phase with $\alpha=0.5$, $\beta=1.5$, and $\mu=1$, where $\sigma^{\rho}_{i}$ scales as $\mathcal{O}(1/L)$, thereby vanishing in the thermodynamic limit.
Figure 8: Plots of the relative fluctuation, $\sigma_{a}/\langle N_{a}\rangle$, of the reservoir populations $N_{a}$ ($a=1,2$ for reservoirs $R_{1}$ and $R_{2}$) as a function of system size $L$. $\sigma_{a}=\sqrt{\langle N_{a}^{2}\rangle-\langle N_{a}\rangle^{2}}$ is the standard deviation of $N_{a}$, and $\langle.\rangle$ denotes the time-average. (a) DW–DW phase with complete delocalization for $\alpha=\beta=0.5$ and $\mu=1$. (b) LD–LD phase for $\alpha=0.5$, $\beta=1.5$, and $\mu=1$. In both cases, the relative fluctuation decreases monotonically with increasing $L$.
@@ VI Phase boundaries and phase transitions
The phase diagrams of our model (Figs. 2 and 10) differ substantially from that of the open TASEP, or from other models of TASEPs with finite resources. We identify the boundaries delineating different phases according to MFT. This can be achieved by equating the currents in corresponding phases.
@@
Boundaries (72) and (73) appear as straight lines parallel to the $\beta$ and $\alpha$ axes, respectively, see the phase diagrams in Figs. 2(b)-2(d) and Figs. 10(a)-10(b). The requirement $\alpha,\beta>0$ in Eqs. (72) and (73) imposes the following range on $\mu$ for which these boundaries can exist:
@@
Depending on the value of $\mu$, these boundaries can be either straight lines or curved ones, as shown in Figs. 2 and 10.
We now discuss the nature of the phase transitions across these phase boundaries. The $\alpha$-$\beta$ phase space of our model can exhibit the presence of either two or four phases in different regions of the $\alpha$-$\beta$ plane, depending sensitively on the specific value of $\mu,\;0\leq\mu\leq 2$. For $0\leq\mu\leq 1/2$, only the LD-LD and DW-DW phases are present; while for $3/2\leq\mu\leq 2$, only the HD-HD and DW-DW phases are observed. In contrast, for intermediate values ($1/2\leq\mu\leq 3/2$), all four phases – LD-LD, HD-HD, MC-MC, and DW-DW – are observed in different regions of the $\alpha$-$\beta$ plane, the regions meeting at a common point in the $\alpha$-$\beta$ plane. Due to particle-hole symmetry described by (10) - (11), the phase diagram for $\mu>1$ can be obtained from that for $\mu<1$. Furthermore, considering the bulk densities as the order parameters, the transitions between different phases – for example, from LD-LD or HD-HD to MC-MC or DW-DW – are continuous transitions, since the bulk density does not show any jump across any of the phase boundaries. These are summarized in Table 2.
Interestingly, Figs. 2 and 10 reveal that, for $1/2<\mu<3/2$, all four phase boundaries intersect at a common point. Since the transitions across the different phase boundaries are all continuous transitions, the transitions across this common meeting point are also continuous in nature. We call this a multicritical point; see Refs. [3, 4, 5, 10] in this context, and Ref. [64] for discussions on multicritical points in equilibrium systems. The coordinates of this multicritical point in the $\alpha$-$\beta$ phase diagram can be obtained by simultaneously solving the equations (72), (73), (75), and (76):
@@
As $\mu$ varies, the multicritical point ($7$) traces a hyperbolic trajectory [see Fig. 9(a)]:
@@
Notably, $d$ diverges as $\mu\rightarrow(1/2)^{+}$ and $\mu\rightarrow(3/2)^{-}$, indicating that the multicritical point exists only within the range $1/2<\mu<3/2$ [see Fig. 9(b)]. The symmetric nature of the plot of $d$ versus $\mu$ about $\mu=1$ reflects the particle-hole symmetry of the model.